If the sides of a square are decreased by 2 cm, the area is decreased by 36cm^2. What were the dimensions of the original square?

Answer by checkley75(3666)

X^2=AREA OF ORIGINAL SQUARE
X^2-36=(X-2)(X-2)
X^2-36=(X-2)(X-2)
X^2-36=X^2-4X=4
X^2-X^2+4X=4+36
4X=40
X=40/4
X=10 ANSWER FOR THE ORIGINAL SIDE LENGTH.
PROOF
10*10-36=(8*8)
100-36=64
64=64